### v1

# Robust Deep Network Learning of Nonlinear Regression Tasks by Parametric Leaky Exponential Linear Units (LELUs) and a Diffusion Metric

Abstract
This document proposes a parametric activation function (ac.f.) aimed at improving multidimensional nonlinear data regression. It is a established knowledge that nonlinear ac.f.’s are required for learning nonlinear datasets. This work shows that smoothness and gradient properties of the ac.f. further impact the performance of large neural networks in terms of overfitting and sensitivity to model parameters. Smooth but vanishing-gradient ac.f.’s such as ELU or SiLU have limited performance and non-smooth ac.f.’s such as RELU and Leaky-RELU further impart discontinuity in the trained model. Improved performance is demonstrated with a smooth “Leaky Exponential Linear Unit”, with non-zero gradient that can be trained. A novel diffusion-loss metric is also proposed to gauge the performance of the trained models in terms of overfitting.
keywords:
Deep Neural Network , Nonlinear Regression , Overfitting , Activation Function , Diffusion Loss
[…]
organization=Independent Researcher, addressline=Starnberger Weg 7, city=Gilching, postcode=82205, country=Germany
1 Introduction
It is typical in engineering applications that a modelled system is multidimensional, i.e., a representative output parameter of the system depends on many system operational variables. The model might also be structured following a mesh-like combination of the system variables, which were input on a computational model or test bench for the data generation. In the example to be addressed in this document, a metric of an electric motor efficiency is dependent on three variables, representative of the motor speed, power, and electric feed. This seemingly simple problem has proven to be surprisingly difficult to be trained for data regression. In an effort on Gaussian Process Regression (GPR) methods [1], the author developed a loss based on a diffusion operator to detect and avoid overfitting resulting from the highly nonlinear training landscape.
An argument in [2] stating that deep-learning methods naturally provides a Occam’s Razor behaviour has instigated the author to apply a neural network (NN) to the problem. There was an expectation that this property would naturally avoid overfitting as seen with the GPR models. The nature of the motor model has also raised some learnings in this deep learning context and are herein reported.
The scientific and engineering communities are interested in modelling techniques that are robust to parameters and that provide good generalisation. The selection of the activation function (ac.f.) of a NN heavily impacts such objectives, and it is known that nonlinear ac.f.’s are necessary to be able to model nonlinear datasets [3]. However, most of the literature focuses on large models for classification problems [4, 5, 6, 7], and regression problems are typically small [8] and the concern is usually with underfitting.
A dense feedforward NN was set up to learn the aforementioned motor model, for which relatively large depths and widths were required. In general, the “global” overfitting patterns of the GPR methods seen in [1] are not observed with the NN. As shown in Sect. 4.2, rather “localised” low-amplitude small-lengthscale features are observed – in this sense, “localised” refers to over-expression between adjacent data points, thus not a “global” long wavelength pattern spanning through many data points. In special, the regions where two of the variables are fast changing but the third one is not, seemed to exacerbate the issue.
A motivation to pursue a deeper understanding of the observed issues comes from this remark in [9], that “a priori knowledge of a particular task into a learning algorithm (…) generally improves performance, if the incorporated knowledge is relevant to the task.” The author proposes an ad hoc metric of the flexibility of the ac.f. as a ratio of its maximum and minimum gradients. In considering this metric, it can be observed that highly flexible ac.f.’s are prone to over-expressing in large models, whereas the conventional linear unit has zero flexibility and thus only represents linear functions (a common knowledge [3]). The current work proposes a parametric, trainable flexibility that enhances the a priori “fitness” of an ac.f. for more generalisable nonlinear learnings.
Furthermore, concern with ac.f. smoothness has motivated other work [5, 7], and the notion of trainable/adaptive coding in ac.f. has demonstrated successful behaviour [10, 11, 12]. This outcome is herein achieved with a simple ac.f. definition, termed Leaky Exponential Linear Unit (Leaky ELU, or LELU). Importantly, the focus here is broader in the sense that it focuses on a set of behavioural features of an ac.f. for generalisation, one of them being the discussed flexibility, but also robustness (i.e., reduced dependency) with model size. Quoting [12] “there is no rule of thumb of choosing an optimal activation function.” Essentially, we want to decrease the effort in finding an optimal ac.f. by providing an adaptive framework that is able to code the degree of presence of particular phenomena in the input in a smooth and ever proportional way.
Another proposition in the current work is the use of a diffusion-like metric to gauge the performance of the NN, especially in terms of overfitting. As developed in [1], the “entropy-in-features” of the training data is computed via a modified Laplace operator and considered as a true label. Testing performed on a staggered mesh measures the entropy content of a trained model with respect to the true label, essentially meaning that a trained model should limit the amount of extra features it determines, and thus less “noise” stemming from overfitting.
The localised overfitting behaviour is demonstrated on canonical unidimensional problems, with model sizes representative of that for the motor model. It is shown that classical ac.f.’s have mixed performance in the problems whereas the proposed LELU is robust to model size. The observed properties of the LELU are further demonstrated in the motor model.
2 Activation Functions
The author herein proposes an ad hoc score of the “flexibility” (term inspired by [13]) of an ac.f. to serve as a proxy of its behaviour in data regression tasks. The flexibility score $\eta(\phi)$ is proportional to the ratio of the maximum and minimum slopes of an ac.f. $\phi$, such that
$\eta(\phi)=1-\frac{\min(\phi^{\prime})}{\max(\phi^{\prime})}.$ (1)
This scores is zero for inflexible ac.f.’s, such as the Linear Unit (LU), which can only represent linear functions. On the other extreme, an ac.f. with score 1 is highly flexible. Table 1 shows the score for the ac.f.’s plotted in Fig. 1.
Table 1: Flexibility score of classical activation functions by application of Eq. 1. The parametrized activation functions affect the flexibility score by their control parameter.
Activation $\eta(\phi)$
LU 0
tanh 1
ReLU 1
ELU [5] 1
Leaky ReLU [14, 4] $1-\alpha$
SiLU [15] 1.1
Softplus [9] 1
LELU $1-\beta$
Figure 1: Graph of several activation functions.
In the current work, it is shown that for nonlinear regressions tasks the following features of an ac.f.’s negatively impact the model performance:
  • 1.
    a higher flexibility score is more likely to present overfitting,
  • 2.
    discontinuous functions impart discontinuity into the trained data,
  • 3.
    vanishing gradient impairs training, irrespective if the function saturates at non-zero values [14],
  • 4.
    absence of negative values is too restrictive for neuron activation for smooth functions and limits their learning.
This motivates the proposal of a parametric ac.f. with a trainable parameter that overcomes these limitations and thus allows for faster and more robust training for nonlinear regression tasks. The family of activation functions, termed “Leaky Exponential Linear Unit” (Leaky ELU, or LELU), is given by
$\text{LELU}(x,\beta)=\begin{cases}~{}x&\text{if }x>0\\ ~{}\exp((1-\beta)x)-1+\beta~{}x&\text{if }x\leq 0\end{cases}$ (2)
where the trainable parameter $\beta$ controls the slope of the activation function in the negative region (see Fig. 1).
The parametric LELU is constructed by imposing the following objectives for its features and behaviours
  1. 1.
    monotonically continuous,
  2. 2.
    non saturating (non-null gradients),
  3. 3.
    to push the unit mean activation closer to zero.
These requirements are imposed onto the ac.f. construction to enhance the performance for nonlinear regressions tasks. The regression results show that this set of features inherently produces a behaviour similar to batch normalisation without the need for regularisation, as for instance explicitly required in [7, 16].
3 Model Description
The general aspects of the deep-learning model are initially presented. In the sequence, the diffusion-loss metric formulation is described.
3.1 Neural Network Configuration
A fully-connected feedforward NN is set up with $n$ hidden layers each containing $m$ neurons, and a linear-uniform single-neuron output, as sketched in Fig. 2.
Figure 2: A vanilla feedforward neural network of depth $n$ and width $m$, with three inputs and one output. The output layer is linear.
The NN is implemented in Python using the Keras library [17] and trained to minimise the mean absolute error (MAE) employing the Adam optimizer [18]. The minimisation of the MAE has provided better results for the regression tasks when compared to a mean squared error (MSE). The learning rate is step-annealed based on the MAE loss – more details are later provided for each model. Neurons are initialised with a HeNormal distribution [4].
Regularization in regression tasks is essentially a trade-off: it helps prevent overfitting and can improve training stability, but by definition the minimum achievable training error is limited by the strength of the regularization. Therefore, no regularisation is used as default in the current work, nonetheless this aspect is also herein briefly assessed since it is a common practice.
The use of a custom diffusion operator as a testing metric of performance allows for training on the full dataset. This metric is assessed via its MSE and described in the following subsection.
3.2 Diffusion-Loss Metric
The diffusion-loss formulation herein applied follows the more complete description in [1]. This methodology allows for the creation of a hierarchical testing metric from the training data field without requiring splitting a training dataset into testing sets. Training is performed on the complete available dataset and offers best chance for generalisation.
The method for identifying oscillations uses a measure of the diffusion of the training data represented by a modified Laplace operator $\nabla$. This sensor in a 1d finite-difference context is computed as follows
$\nabla y_{i}=\frac{1}{\Delta^{2}}\frac{\big{|}y_{i+1}-2y_{i}+y_{i-1}\big{|}}{y_{i+1}+2y_{i}+y_{i-1}},$ (3)
for the $i$-th point of the mesh in Fig. 3, with $\Delta$ representing the mesh spacing, and $y$ a positive property. In the currently proposed method, this diffusion sensor applied to the training data serves as the true label.
Figure 3: A one-dimensional mesh: the function values are shown by squares and the mesh nodes by circles. The original training mesh is shown as black nodes and the created staggered nodes, in red with half indices. Brown and orange squares show possible oscillations around node $i$.
In order to create a suitable testing metric sensitive to overfitting in between training points, staggered nodes centred between the training points are created (shown in red in Fig. 3). An updated diffusion sensor $\tilde{\nabla}$ applied to these staggered nodes, indexed by half indices, is proposed
~$\tilde{\nabla}y_{i}=\frac{1}{3~{}(\Delta/2)^{2}}\frac{\big{|}\hat{y}_{i+1}-\hat{y}_{i+\frac{1}{2}}-\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}\big{|}}{\hat{y}_{i+1}+\hat{y}_{i+\frac{1}{2}}+\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}},$ (4)
with $\hat{y}$ representing predicted function values, $\hat{y}_{i}=f(x)_{i}$, with $\hat{y}=y$ expected at the training points. This stencil is able to identify the “ringing/wiggling” features around the central node $i$. Therefore, deviations between a testing sensor (Eq. 4) and the true score (Eq. 3) can thus be used as testing metric in model training.
The creation of a staggered mesh in one dimension as shown in Fig. 3 is straightforward since a central node only has a single neighbour to either its side. On a multidimensional case, an approach based on a single staggered mesh built on centroids of the original mesh cells is proposed. In this approach, we are required to use the centre-crossing diagonals (composed of pairs of antipodal points) as an ad-hoc means to compute Laplace operators. The resulting staggered mesh for a 3d case is shown in Fig. 4.
Figure 4: The original multidimensional training mesh shown as black points and the created cell-centred staggered points shown in red, with half indices. The diffusion sensors are computed along the centre-crossing diagonals shown in red. Filled symbols are used for the sensor computations for the central node ($(i,j,k)$ in 3d, for instance).
The staggered Laplace-operator diffusion sensor written for a centre-crossing diagonal $p$ of the node $(i,j,k)$ in Fig. 4 is given by
~$\tilde{\nabla}y_{i,j,k}^{p}=\frac{1}{3~{}(\Delta/2)^{2}}\frac{\big{|}y_{i+1,j+1,k+1}-y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}-y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}\big{|}}{y_{i+1,j+1,k+1}+y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}+y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}},$ (5)
and for the true label
$\nabla y_{i,j,k}^{p}=\frac{1}{\Delta^{2}}\frac{\big{|}y_{i+1,j+1,k+1}-2y_{i,j,k}+y_{i-1,j-1,k-1}\big{|}}{y_{i+1,j+1,k+1}+2y_{i,j,k}+y_{i-1,j-1,k-1}}.$ (6)
The complete 3d diffusion sensors for the node $(i,j,k)$ are given by the summation over all centre-crossing diagonals resulting in
$\nabla y_{i,j,k}=\sum_{p=1}^{4}\nabla y_{i,j,k}^{p}~{},~{}~{}~{}\tilde{\nabla}y_{i,j,k}=\sum_{p=1}^{4}\tilde{\nabla}y_{i,j,k}^{p}.$ (7)
Note that the diffusion sensors are only computed at the interior of the original mesh, where the complete stencil is available. The reader is referred to [1] for further and more general discussions and information about the methodology as well as snippets of code implementation.
4 Results and Discussions
We are interested in multidimensional problems that typically require large models to capture all their features. The author argues that using canonical problems with isolated features and fitting small NNs to them do not guarantee generalisation of the approach for those large, real-world problems. We want to be able code the degree of presence of series of particular phenomena that might compose large problems. Therefore, in this work, we take large NNs that are required to represent large, complex models, and apply them to canonical problems in search of ac.f.’s that offer robustness whilst coding the degree of presence of different and composed phenomena. The goal is to provide an ac.f. that works well regardless of model size and that can scale effectively from small to large networks, thereby enabling potentially more robust and less time consuming regression tasks.
4.1 Unidimensional Problems
The proposed unidimensional problems [^1] Source files available in http://github.com/ebiga/nn_1d_overfit_ex represent certain important features of the target motor model. A hyperbolic tangent function is chosen for its non-monotonic change of slope, whereas an exponential function is chosen for its fast change of properties. Numerical experiments with a NN of depths $n=[7,8]$ and widths $m=[120,240]$ (fixed per layer) are executed for the ac.f.’s listed in Table 2.
4.1.1 Hyperbolic Tangent Function
A 1d equation given by $f(x)=0.5+0.5\tanh(5x),~{}x\in[-1,1]$ is used to generate two sets of equally spaced, noiseless $(7,14)$ training points. The model is trained for 15000 epochs with batch size 3. The learning rate is scheduled in the range $[1e-3,1e-6]$ with factor 0.5, patience 500, and cooldown 100.
Table 2: Losses of the trained models for the hyperbolic tangent data.
Activation Training MAE ×$\times 10^{-6}$ Diffusion MSE ×$\times 10^{-3}$
$14pts$ $7pts$ $14pts$ $7pts$
Leaky ReLU $(\alpha=0.2)$ 14 10 90 216
LELU $(\beta=0.3)$ 20 27 16 25
LELU $(\beta=0.4)$ 22 101 49 72
LELU $(\beta=0.6)$ 54 68 15 73
Softplus 71 21 16 220
SiLU 23 19 13 49
tanh 17 8 156 372
ELU 17 88 55 234
ReLU 8 5 125 241
Table 2 presents the training losses and diffusion metrics of the trained model. The MAEs are generally low and care must be taken that the models are not tending to memorise the training points, the performance of the models is therefore truly assessed by the diffusion metric. The SiLU and LELU ($\beta=0.3$) achieve low MAE and low diffusion losses, and the performance is not strongly lost with the reduction of the number of training points. This result is indicative of less overfitting, which is corroborated by the function plots in Fig. 5. Strongly overfitting is observed with some ac.f.’s in this setup, whereas the proposed LELU is relatively robust to the change in the number of training points. Not shown for the sake of conciseness, very similar results are observed with a $7\times 120$ model as well.
(a) 7 training points.
(b) 14 training points.
Figure 5: Trained hyperbolic tangent function plots for the different activation functions.
4.1.2 Exponential Function
A 1d exponential equation given by $f(x)=0.1^{x},~{}x\in[-1,1]$ is used to generate a set of $12$ equally spaced, noiseless training data points. The model is trained for 20000 epochs with batch size 3. The learning rate is scheduled in the range $[1e-3,2e-7]$ with factor 0.5, patience 500, and cooldown 100.
Table 3: Losses of the trained models for the exponential function data.
Activation Training MAE ×$\times 10^{-6}$ Diffusion MSE ×$\times 10^{-3}$
$n=8$ $n=7$ $n=8$ $n=7$
Leaky ReLU $(\alpha=0.2)$ 4 31 12 16
LELU $(\beta=0.3)$ 50 19 3 2
LELU $(\beta=0.4)$ 24 33 2 2
LELU $(\beta=0.6)$ 34 62 1 1
Softplus 120 133 5 2
SiLU 7 25 13 14
tanh 21 14 28 37
ELU 18 12 18 20
ReLU 4 5 25 19
Table 3 presents the training losses and diffusion metrics of the trained model. As in the previous case, the lower the MAE, the higher is the tendency of the model to memorise the training data, with a resulting increase in the diffusion metric. For this case, both SiLU and Softplus show a poorer performance and trade the training and diffusion (overfitting) errors. The LELUs are able to adequately meet both low MAE and diffusion errors. The trained model predictions are plotted in Fig. 6. As in the previous case, different levels of overfitting are observed with the classical ac.f.’s in this setup, whereas generally robust results are obtained with the LELU.
(a) Depth 7.
(b) Depth 8.
Figure 6: Trained exponential function plots for the different activation functions.
4.1.3 Exponential Function and a Single Neuron
The previous problem is reduced to a set of three training points, $x_{i}=(-1,0,1)$, which a single neuron shall learn. This exercise can be performed by hand and two of the three points can be exactly matched. The obtained neuron parameters are used to infer a function as shown in Fig. 7, plotted along the ac.f. derivatives and training errors.
Figure 7: Predictions with a single neuron trained for the three data points.
In terms of training error (at the third point), ReLU, SiLU and Softplus produce smaller errors whereas the LELU has the largest. However, ReLU, ELU, SiLU and Softplus tend to null gradients whereas the Leaky ReLU and LELU produce a strong signal that indicates that that behaviour cannot be coded – the task therefore requires and should be taken by the next of kin on a larger model.
From the perspective of learning nonlinear regression tasks, such weak signal is less then ideal. The gradient and Hessian of the loss can be computed by hand for this simple problem. This operation shows that the weak ac.f.’s produce a vanishing loss gradient and higher conditioned Hessians, which is indicative of the tendency to poor learning and overfitting. The LELU on the other hand produces a clear gradient and a lower conditioned Hessian. With this simple exercise, the author intends to anecdotally show that the higher “flexibility” of the ac.f. is incurring in a weak signal for subsequent neurons and is a potential cause of instability showing as overfitting.
4.1.4 Other Findings with the Unidimensional problems
(a) Hyperbolic tangent, with 7 points.
(b) Exponential function, depth 8.
Figure 8: Trained function plots for the different activation functions on a positive range of $x$.
It is interesting to note that some ac.f.’s struggle around $x=0$ in both case studies, even when this point is not a training data. A thorough explanation is challenging for a large model [16], but the author believes this can be partially understood by a similar analysis of the gradient as in the previous subsection. The loss-gradient is proportional to the summation of the gradient of the ac.f. times the training point coordinate. If the training point coordinate tends to zero, or adjacent training points cross $x=0$, this can further induce the loss gradient summation to vanish or “flicker”, and to destabilize the learning on an already unstable model setup. This behaviour is confirmed with a simple shift in $x$ by two units, as shown in Fig. 8, in which the local instability around $x=0$ is removed for the unstable models. Note that SiLU actually lost performance and presents overfitting, opposed to the behaviour shown in Fig. 5a, which corroborates the borderline stability of the model.
Another observation is that discontinuous ac.f.’s generate discontinuous trained models in both exercises. In these and some other setups not shown in here, the discontinuous ac.f.’s rather often default to piecewise linear connection between the training points. Even when this was not the case, the trained model is still discontinuous, as seen in Figs. 5, 6 and 8.
The adopted canonical problems indicate that classical ac.f.’s lack robustness with the problem setup for nonlinear regressions tasks when trained by large models. The proposition of the objectives in Sect. 2 for the LELU construction aims at mitigating the reported issues.
4.1.5 Other Approaches to Decrease Overfitting
By assuming that the author has been successful in the beginning of this section to justify the most obvious criticism against these exercises, i.e., the adopted model sizes, other strategies also merit further discussions. Known strategies to deal with overfitting, namely, regularisation and smaller learning rates for longer epochs were also evaluated. The results are not shown for the sake of conciseness since the outcomes are as expected, i.e., a decrease in the overfitting is achieved however at the expense of considerably higher training MAEs and/or computational resources. Both L1 and L2 regularisations, with constant strength $41e-$ 4, are able to reduce or remove (L1 regularisation) overfitting, however at the expense of $10x$ or larger training MAEs. In terms of learning rates, an example with the ELU required reduction of half of the initial learning rate, and $10x$ longer training epochs.
The LELU is able to achieve adequate results without those penalties. It can robustly learn the nonlinear regression task without memorising the training data. This behaviour is further confirmed on the more complex model discussed in what follows.
4.2 Electric Motor Model
The noise-free system model described in the introduction is composed of a 3d mesh of (param1, param2, param3) = $(19\times 15\times 5)$ points, shown in Fig. 9.
Figure 9: Wireframe view of the motor model mesh with the third dimension shown as the “thickness” of the wireframe. The third dimension collapses at the lower borders.
The problem shows a fast change of behaviour along $param1$ and an even faster one with $param2$, which combine at one corner of the operational envelope of the motor. The behaviour with $param3$ shows as the “thickness” of the wireframe, and it changes with both other parameters and collapses at their lower borders. A nonlinear feature can also be observed approaching the maximum limits of the $param1$ range.
The dataset is normalised so as to keep unitary mesh spacing in all dimensions to avoid cell skewness (which also favours the computation of the diffusion differentials), and kept in the positive range to support the more sensitive ac.f.’s. For the current study the trained data must exactly reproduce the noise-free training information [1]. A noiseless training is generally more challenging and is an additional reason for this choice in the current exercise.
A numerical experiment with the implemented [^2] Source files available in http://github.com/ebiga/gpr_aniso_trials NN with combinations of depth $n=[6,7,8]$ and width $m=[80,120,240,360]$ (fixed per layer) is executed. The model is trained for 9000 epochs with batch size 35. The learning rate is scheduled in the range $[1e-3,1e-5]$ ($45e-$ 4 for the models with width $360$ and the $8\times 240$ case) with factor 0.5, patience 250, and cooldown 50.
(a) MAE.
(b) Diffusion SE.
Figure 10: Trained model losses for the electric motor.
Figure 10 shows the MAE and diffusion metrics for the different model sizes (indexed by $m^{n}/1e12$ in the plots). The tendency to overfit shows as a lack of “convergence” of the MAE with the increase in the model size. For this complex problem the mean errors do not show large differences between the “good” and “bad,” so visual inspection is also necessary. Nonetheless, it can be seen that both metrics are lower with the LELU and are also more robust to model size. SiLU and Softplus do not present any favourable behaviour as partially seen with the canonical problems. The conclusions above are corroborated by the slice plots in Figs. 11 and 12.
(a) LELU $(\beta=0.2)$
(b) LELU $(\beta=0.4)$
(c) ELU
(d) SiLU
(e) Softplus
(f) Leaky ReLU $(\alpha=0.2)$
Figure 11: Trained electric motor model predictions sliced at $(param1,param3)=(28,0.8)$ for different activation functions and model sizes.
Figure 11 shows slices along $param2$, which is the fast-changing variable. The tanh and ReLU results are not shown due to the poor performance. The region $param2=[0,2]$, where the highest slopes are seen, presents a tough task for the ac.f.’s. Considerable “localised” overfitting is observed for the ELU, SiLU and Softplus. Interestingly, the less flexible leaky ac.f.’s show lower overfitting content and also less dependency on the model size. This nice behaviour prompted the current effort into assembling this property into a new ac.f. The Leaky ReLU results are “kinky” though and such behaviour is not ideal. The proposed LELU with $\beta=0.4$ shows a robust performance, indicating that the lower flexibility and smoothness are indeed advantageous.
(a) LELU $(\beta=0.2)$
(b) LELU $(\beta=0.4)$
(c) ELU
(d) SiLU
(e) Softplus
(f) Leaky ReLU $(\alpha=0.2)$
Figure 12: Trained electric motor model predictions sliced at $(param1,param2)=(28,7.2)$ for different activation functions and model sizes.
Figure 12 shows similar slices along the tight- and slow-changing $param3$. This slice allows for interesting behaviours to be more closely assessed, i.e., the spread of the ac.f. with different model sizes and the capacity to code small values. The less flexible leaky ac.f.’s clearly show tighter and more representative capacity. SiLU and Softplus are also able to demonstrate a tighter spread, albeit with an offset error to the training data, whereas ELU is generally poorly behaved.
By inspection of ac.f. graphs in Fig. 1, it can be seen that both SiLU and Softplus present a wider range of slowly changing gradient, or in other words, ELU has a sharper gradient change. In the SiLU case, the gradient is actually never zero in the range plotted in Fig. 1. This observation has further prompted the effort into coding this property into the LELU. As seen in Fig. 1, this ac.f. slowly changes slopes and is therefore smoother than other ac.f.’s, thereby enabling a stronger a priori fitness for generalisable nonlinear regressions.
(a) $6\times 120$
(b) $7\times 120$
(c) $8\times 120$
Figure 13: Model size effect on the electric motor predictions for different activation functions.
For this motor model the LELU with $\beta=0.4$ provides the most agreeable results. Upon careful visual inspection of each NN result (see Fig. 13), depth 6 is not enough to capture all features, whereas depth 7 is able to do so but the MAE still indicates some distance (error) to the training points. Generally at least width 120 is necessary for acceptable results.
Figure 14: Regularisation effect on the electric motor predictions for the $8\times 120$ model.
It is worth revisiting the discussion in Sect. 4.1.5. Except from reducing the model size, which is in fact hopefully clear now the need for such larger models, the other techniques to control overfitting also apply to the motor model. As shown in Fig. 14, both L1 and L2 regularisations (strength $41e-$ 4) reduce the overfitting behaviour of the SiLU, at the expense of representation of training points. The MAE increase by more than a factor of 6 in these cases and the performance does not improve over the LELU ($\beta=0.4$). Another strategy is to implement a change of variables, for instance a power correction $\tilde{y}=y^{0.1}$, to reduce the large differences in the max and min values. A model trained on this function learns faster and easier, however the resulting errors when brought back to the original scale are also very large, which defeats the purpose. As already observed with the canonical problems, the LELU seems to provide adequate results without those drawbacks and extra steps.
5 Concluding Remarks
A parametric activation function for nonlinear regressions tasks that is monotonically smooth, with a non-null gradient at the negative region, is developed. The proposed ac.f. is termed Leaky Exponential Linear Unit (or Leaky ELU, or LELU). The non-null gradient is proposed as a means to control an ad hoc score of the flexibility of an ac.f. and is a trainable parameter. Available nonlinear ac.f.’s have generally high flexibility scores which seem to be associated with proneness to overfitting. A null flexibility score can only represent linear data, on the other hand.
Smoothness is sought to avoid coding discontinuities at the predictions of the trained model. This is shown to be a result of the ReLU and Leaky ReLU ac.f.’s. The Leaky ReLU however has a trainable flexibility and this is shown to support the claim of beneficial and robust behaviour of this property for nonlinear regression tasks.
The results with the LELU are promising for highly nonlinear regressions tasks and large models. Its properties implicitly produce a behaviour similar to regularisation without strong penalisation of the training errors. The LELU does not show evidence of trading off training error with the overfitting/testing metric, at least not as strongly as other ac.f.’s. It also shows less sensitivity with model size and training points for nonlinear regression tasks.
Another proposition in the current work is a metric to identify overfitting. The method is based on a modified Laplace operator and tests for oscillation in the interior of the training mesh cells. A true label is provided by the modified Laplace operator applied to the original training nodes, which acts as a measure of the “entropy-in-features” contained in the training data. A trained model must limit (ideally avoid) the amount of extra features it determines based on its diffusion metric. The testing points must be created on a staggered fashion to be able to gauge any excess entropy in the interior of the mesh cells. With the proposed method, the model can be trained on all available training points.
The author expects the proposed methods to be generalisable to other machine learning or dataset strategies. The observed properties of the proposed LELU should reduce the workload to select a good model setup and is as a such a highly desirable behaviour. The Laplace operator with its measure of the diffusion of a system provides a powerful tool for detection (and prevention) of overfitting in data regression. They should provide a sure hand in the process of implementing a model by bringing more robustness and clearer metrics.

### v2

@@
 This document proposes a parametric activation function (ac.f.) aimed at improving multidimensional nonlinear data regression.
-It is a established knowledge that nonlinear ac.f.’s are required for learning nonlinear datasets.
+It is a established knowledge that nonlinear ac.fs.
+are required for learning nonlinear datasets.
 This work shows that smoothness and gradient properties of the ac.f.
 further impact the performance of large neural networks in terms of overfitting and sensitivity to model parameters.
-Smooth but vanishing-gradient ac.f.’s such as ELU or SiLU have limited performance and non-smooth ac.f.’s such as RELU and Leaky-RELU further impart discontinuity in the trained model.
+Smooth but vanishing-gradient ac.fs.
+such as ELU or SiLU (Swish) have limited performance and non-smooth ac.fs.
+such as RELU and Leaky-RELU further impart discontinuity in the trained model.
 Improved performance is demonstrated with a smooth “Leaky Exponential Linear Unit”, with non-zero gradient that can be trained.
@@
 Deep Neural Network , Nonlinear Regression , Overfitting , Activation Function , Diffusion Loss
+journal: Information Sciences
 […]
@@
 1 Introduction
+1.1 Perspectives and Motivation
 It is typical in engineering applications that a modelled system is multidimensional, i.e., a representative output parameter of the system depends on many system operational variables.
-The model might also be structured following a mesh-like combination of the system variables, which were input on a computational model or test bench for the data generation.
+The model might also be structured following a mesh-like combination of the system variables, which were input on a computational model or a test bench for the data generation.
 In the example to be addressed in this document, a metric of an electric motor efficiency is dependent on three variables, representative of the motor speed, power, and electric feed.
 This seemingly simple problem has proven to be surprisingly difficult to be trained for data regression.
-In an effort on Gaussian Process Regression (GPR) methods [1], the author developed a loss based on a diffusion operator to detect and avoid overfitting resulting from the highly nonlinear training landscape.
+In an effort with Gaussian Process Regression (GPR) [1], the author developed a loss based on a diffusion operator to detect and avoid overfitting resulting from the highly nonlinear training landscape.
 An argument in [2] stating that deep-learning methods naturally provides a Occam’s Razor behaviour has instigated the author to apply a neural network (NN) to the problem.
@@
 The scientific and engineering communities are interested in modelling techniques that are robust to parameters and that provide good generalisation.
-The selection of the activation function (ac.f.) of a NN heavily impacts such objectives, and it is known that nonlinear ac.f.’s are necessary to be able to model nonlinear datasets [3].
-However, most of the literature focuses on large models for classification problems [4, 5, 6, 7], and regression problems are typically small [8] and the concern is usually with underfitting.
-A dense feedforward NN was set up to learn the aforementioned motor model, for which relatively large depths and widths were required.
-In general, the “global” overfitting patterns of the GPR methods seen in [1] are not observed with the NN.
+The selection of the activation function (ac.f.) of a NN heavily impacts such objectives, and it is known that nonlinear ac.fs.
+are necessary to model nonlinear datasets [3].
+Furthermore, smoothness (e.g., the classical ELU [4]) as well as trainable/adaptive coding properties for ac.fs.
+receive consistent attention in the literature, and the reader is referred to the thorough summary in [5] for further information.
+More complex approaches such as ac.fs.
+based on GPRs [6], rational functions [7], Kronecker Neural Networks [8], or even on model structural changes [9], have also been proposed.
+Of importance to the current work, several trainable variants are available solely for ReLU and ELU (again summarised in [5]).
+The choice of an ac.f.
+can in fact be overwhelming and other broad-scope reviews are available in [10, 11].
+However, the literature generally focuses on large models for classification problems whereas regression problems are typically small.
+Even though such reviews provide interesting background for further exploration, regression tasks can be more demanding than classification since they require continuity.
+An interesting example of such difficulties is a compromise termed “benign overfitting” [12], in which localised spikes near noisy training points are accepted for a model that generalises well elsewhere.
+In the current work, a dense feedforward NN is developed to learn the aforementioned motor model.
+Relatively large depths and widths are required to properly learn such nonlinear problem [13, 14].
 As shown in Sect.
-4.2, rather “localised” low-amplitude small-lengthscale features are observed – in this sense, “localised” refers to over-expression between adjacent data points, thus not a “global” long wavelength pattern spanning through many data points.
+4.2, over-expression between adjacent data points is observed with classical ac.fs.
+and the use of regularisation to avoid such patterns incur in larger training errors.
 In special, the regions where two of the variables are fast changing but the third one is not, seemed to exacerbate the issue.
@@
 as a ratio of its maximum and minimum gradients.
-In considering this metric, it can be observed that highly flexible ac.f.’s are prone to over-expressing in large models, whereas the conventional linear unit has zero flexibility and thus only represents linear functions (a common knowledge [3]).
+In considering this metric, it can be observed that highly flexible ac.fs.
+are prone to over-expressing in large models, whereas the conventional linear unit has zero flexibility and thus only represents linear functions (a common knowledge [3]).
+1.2 Contributions of the Work
 The current work proposes a parametric, trainable flexibility that enhances the a priori “fitness” of an ac.f.
 for more generalisable nonlinear learnings.
-Furthermore, concern with ac.f.
-smoothness has motivated other work [5, 7], and the notion of trainable/adaptive coding in ac.f.
-has demonstrated successful behaviour [10, 11, 12].
 This outcome is herein achieved with a simple ac.f.
@@
 Testing performed on a staggered mesh measures the entropy content of a trained model with respect to the true label, essentially meaning that a trained model should limit the amount of extra features it determines, and thus less “noise” stemming from overfitting.
+1.3 Organisation
+The rationale for the ac.f.
+definitions are initially discussed.
+The NN setup and diffusion loss definitions are then presented.
 The localised overfitting behaviour is demonstrated on canonical unidimensional problems, with model sizes representative of that for the motor model.
-It is shown that classical ac.f.’s have mixed performance in the problems whereas the proposed LELU is robust to model size.
+It is shown that classical ac.fs.
+have mixed performance in the problems whereas the proposed LELU is robust to model size.
 The observed properties of the LELU are further demonstrated in the motor model.
@@
 $\eta(\phi)=1-\frac{\min(\phi^{\prime})}{\max(\phi^{\prime})}.$ (1)
-This scores is zero for inflexible ac.f.’s, such as the Linear Unit (LU), which can only represent linear functions.
+This scores is zero for inflexible ac.fs., such as the Linear Unit (LU), which can only represent linear functions.
 On the other extreme, an ac.f.
 with score 1 is highly flexible.
-Table 1 shows the score for the ac.f.’s plotted in Fig.
-1.
-Table 1: Flexibility score of classical activation functions by application of Eq. 1. The parametrized activation functions affect the flexibility score by their control parameter.
-Activation $\eta(\phi)$
+Table 1 shows the score for the ac.fs.
+plotted in Fig.
+1, and includes a wide range of original to more modern ac.fs.
+Table 1: Flexibility score of several activation functions by application of Eq. 1. The parametrizable activation functions can affect their flexibility score (examples in Fig. 1).
+Activation $\mathbf{\phi}$ $\eta(\phi)$
 LU 0
@@
 Leaky ReLU [17, 18] $1-\alpha$
-SiLU [15] 1.1
+SiLU* [19] (also, Swish) 1.1
 Softplus [15] 1
 LELU $1-\beta$
+  • 1.
+    *The term SiLU instead of Swish [20] is adopted for consistency with the Keras library. For the current purposes, the other modern option Mish [21] showed very similar performance to SiLU/Swish.
 Figure 1: Graph of several activation functions.
-In the current work, it is shown that for nonlinear regressions tasks the following features of an ac.f.’s negatively impact the model performance:
+In the current work, it is shown that for nonlinear regressions tasks the following features of an ac.f. negatively impact the model performance:
   • 1.
@@
 $\text{LELU}(x,\beta)=\begin{cases}~x&\text{if }x>0\\ ~\exp((1-\beta)x)-1+\beta~x&\text{if }x\leq 0\end{cases}$ (2)
-where the trainable parameter $\beta$ controls the slope of the activation function in the negative region (see Fig. 1).
+where the trainable parameter $\beta$ controls the slope of the activation function in the negative region (see Fig.
+1).
+Interestingly, the elastic ELU in [22] proposes a change in the slope of the ac.f.
+however on the positive, right-hand side.
 The parametric LELU is constructed by imposing the following objectives for its features and behaviours
@@
 construction to enhance the performance for nonlinear regressions tasks.
-The regression results show that this set of features inherently produces a behaviour similar to batch normalisation without the need for regularisation, as for instance explicitly required in [7, 16].
+The regression results show that this set of features inherently produces a behaviour similar to batch normalisation without the need for regularisation, as for instance explicitly remarked in [23, 7, 24].
 3 Model Description
@@
 Figure 2: A vanilla feedforward neural network of depth $n$ and width $m$, with three inputs and one output. The output layer is linear.
-The NN is implemented in Python using the Keras library [17] and trained to minimise the mean absolute error (MAE) employing the Adam optimizer [18].
+The NN is implemented in Python using the Keras library [25] and trained to minimise the mean absolute error (MAE) employing the Adam optimizer [26], as summarised in Algorithm 1.
+Algorithm 1 Training procedure of the artificial neural network.
+1:Training data $(\vec{x},f(\vec{x}))$. $\triangleright$ From system models or generated from canonical functions.
+2:Trained neural network and diffusion-loss metric.
+3:procedure NN Setup
+4:  Define the NN architecture (depth and width, see Fig. 2).
+5:  Configure learning rate annealing and number of epochs for the Adam optimizer.
+6:  Set activation functions. $\triangleright$ Regularisation applied when intended.
+7:  Initialize weights with chosen initializer.
+8:end procedure
+9:while max epoch not reached do
+[…]
+11:end while
+12:procedure Diffusion metric $\triangleright$ Detailed in Sect. 3.2.
+13:  Make predictions at staggered mesh nodes and at training points.
+14:  Compute diffusion terms at the training points according to Eq. 7.
+15:  Compute the diffusion loss applying Eq. 8.
+16:end procedure
 The minimisation of the MAE has provided better results for the regression tasks when compared to a mean squared error (MSE).
-The learning rate is step-annealed based on the MAE loss – more details are later provided for each model.
-Neurons are initialised with a HeNormal distribution [4].
+The learning rate is step-annealed based on the MAE loss and the weights are initialised by proper techniques, with more details later provided for each training case.
 Regularization in regression tasks is essentially a trade-off: it helps prevent overfitting and can improve training stability, but by definition the minimum achievable training error is limited by the strength of the regularization.
-Therefore, no regularisation is used as default in the current work, nonetheless this aspect is also herein briefly assessed since it is a common practice.
-The use of a custom diffusion operator as a testing metric of performance allows for training on the full dataset. This metric is assessed via its MSE and described in the following subsection.
+Therefore, no regularisation is used as default in the current work, nonetheless this aspect is also herein assessed since it is a common practice.
 3.2 Diffusion-Loss Metric
+The use of the diffusion loss in Algorithm 1 provides an independent testing metric of performance and also allows for training on the full dataset, as discussed in the current subsection.
 The diffusion-loss formulation herein applied follows the more complete description in [1].
@@
 An updated diffusion sensor $\tilde{\nabla}$ applied to these staggered nodes, indexed by half indices, is proposed
-~$\tilde{\nabla}y_{i}=\frac{1}{3~{}(\Delta/2)^{2}}\frac{\big{|}\hat{y}_{i+1}-\hat{y}_{i+\frac{1}{2}}-\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}\big{|}}{\hat{y}_{i+1}+\hat{y}_{i+\frac{1}{2}}+\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}},$ (4)
+$\tilde{\nabla}y_{i}=\frac{1}{3~(\Delta/2)^{2}}\frac{\big|\hat{y}_{i+1}-\hat{y}_{i+\frac{1}{2}}-\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}\big|}{\hat{y}_{i+1}+\hat{y}_{i+\frac{1}{2}}+\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}},$ (4)
 with $\hat{y}$ representing predicted function values, $\hat{y}_{i}=f(x)_{i}$, with $\hat{y}=y$ expected at the training points.
@@
 The staggered Laplace-operator diffusion sensor written for a centre-crossing diagonal $p$ of the node $(i,j,k)$ in Fig. 4 is given by
-~$\tilde{\nabla}y_{i,j,k}^{p}=\frac{1}{3~{}(\Delta/2)^{2}}\frac{\big{|}y_{i+1,j+1,k+1}-y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}-y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}\big{|}}{y_{i+1,j+1,k+1}+y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}+y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}},$ (5)
+$\tilde{\nabla}y_{i,j,k}^{p}=\frac{1}{3~(\Delta/2)^{2}}\frac{\big|y_{i+1,j+1,k+1}-y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}-y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}\big|}{y_{i+1,j+1,k+1}+y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}+y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}},$ (5)
 and for the true label
@@
 $\nabla y_{i,j,k}=\sum_{p=1}^{4}\nabla y_{i,j,k}^{p}~,~~~\tilde{\nabla}y_{i,j,k}=\sum_{p=1}^{4}\tilde{\nabla}y_{i,j,k}^{p}.$ (7)
-Note that the diffusion sensors are only computed at the interior of the original mesh, where the complete stencil is available.
+Note that the diffusion sensors are only computed at the interior of the original mesh, where the complete stencil is available. The loss of the diffusion sensor is finally computed as the MSE by
+$\mathrm{MSE}_{diffusion}=\frac{1}{E}\sum_{e=1}^{E}(\nabla y_{e}-\tilde{\nabla}y_{e})^{2},$ (8)
+where $E$ is the internal nodes of the training mesh.
 The reader is referred to [1] for further and more general discussions and information about the methodology as well as snippets of code implementation.
@@
 The author argues that using canonical problems with isolated features and fitting small NNs to them do not guarantee generalisation of the approach for those large, real-world problems.
-We want to be able code the degree of presence of series of particular phenomena that might compose large problems.
-Therefore, in this work, we take large NNs that are required to represent large, complex models, and apply them to canonical problems in search of ac.f.’s that offer robustness whilst coding the degree of presence of different and composed phenomena.
+We want to be able to code the degree of presence of series of particular phenomena that might compose large problems.
+Therefore, in this work, we take large NNs that are required to represent large, complex models, and apply them to canonical problems in search of ac.fs.
+that offer robustness whilst coding the degree of presence of different and composed phenomena.
+In essence, the context within is models with excess capacity surpassing network size guidelines [27] to better fit regions of high non-linearity [14] and that likely exceed the sample size [13, 28].
 The goal is to provide an ac.f.
@@
 The proposed unidimensional problems [^1] Source files available in http://github.com/ebiga/nn_1d_overfit_ex represent certain important features of the target motor model.
-A hyperbolic tangent function is chosen for its non-monotonic change of slope, whereas an exponential function is chosen for its fast change of properties.
-Numerical experiments with a NN of depths $n=[7,8]$ and widths $m=[120,240]$ (fixed per layer) are executed for the ac.f.’s listed in Table 2.
+A hyperbolic tangent function is chosen for its non-monotonic change of slope, whereas an exponential function is chosen for its large and fast change of properties.
+Numerical experiments with NNs of different depths and widths are executed with the previously presented ac.fs.
+A set 4 seeds of He-Normal [18] and 2 seeds of Glorot-Uniform [29] weight initialisers have been generated and the results are presented in the form of mean and standard deviation.
+Where applicable, experiments with L2 regularisation are also reported.
 4.1.1 Hyperbolic Tangent Function
-A 1d equation given by $f(x)=0.5+0.5\tanh(5x),~{}x\in[-1,1]$ is used to generate two sets of equally spaced, noiseless $(7,14)$ training points.
-The model is trained for 15000 epochs with batch size 3.
-The learning rate is scheduled in the range $[1e-3,1e-6]$ with factor 0.5, patience 500, and cooldown 100.
-Table 2: Losses of the trained models for the hyperbolic tangent data.
-Activation Training MAE ×$\times 10^{-6}$ Diffusion MSE ×$\times 10^{-3}$
-$14pts$ $7pts$ $14pts$ $7pts$
-Leaky ReLU $(\alpha=0.2)$ 14 10 90 216
-LELU $(\beta=0.3)$ 20 27 16 25
-LELU $(\beta=0.4)$ 22 101 49 72
-LELU $(\beta=0.6)$ 54 68 15 73
-Softplus 71 21 16 220
-SiLU 23 19 13 49
-tanh 17 8 156 372
-ELU 17 88 55 234
-ReLU 8 5 125 241
-Table 2 presents the training losses and diffusion metrics of the trained model.
-The MAEs are generally low and care must be taken that the models are not tending to memorise the training points, the performance of the models is therefore truly assessed by the diffusion metric.
-The SiLU and LELU ($\beta=0.3$) achieve low MAE and low diffusion losses, and the performance is not strongly lost with the reduction of the number of training points.
-This result is indicative of less overfitting, which is corroborated by the function plots in Fig.
-5.
-Strongly overfitting is observed with some ac.f.’s in this setup, whereas the proposed LELU is relatively robust to the change in the number of training points.
-Not shown for the sake of conciseness, very similar results are observed with a $7\times 120$ model as well.
+A 1d equation given by $f(x)=0.5+0.5\tanh(5x),~x\in[-1,1]$ is used to generate two noiseless sets of equally spaced $14$- and a more challenging $7$-point cases.
+The models are trained with $(n\times m)=(7\times 120)$ and $(8\times 240)$, for 15000 epochs with batch size 3.
+The learning rate is scheduled in the range [1e3,1e6][1e-3,1e-6] with factor 0.5, patience 500, and cooldown 100.
+Table 2: Mean and standard deviation of the losses of different model setups for the hyperbolic tangent with 1414 training points.
+Activation Training MAE $\times 10^{-6}$ Diffusion MSE $\times 10^{-2}$
+$n=7$ $n=8$ $n=7$ $n=8$
+ELU (15;7) (18;8) (39;14) (80;98)
+Leaky ReLU $(\alpha=0.2)$ (10;3) (15;5) (126;56) (156;59)
+LELU $(\beta=0.3)$ (19;12) (27;33) (16;4) (25;4)
+LELU $(\beta=0.4)$ (19;10) (30;99) (18;3) (22;6)
+LELU $(\beta=0.6)$ (22;34) (63;57) (14;3) (18;5)
+ReLU (7;3) (7;3) (216;59) (275;34)
+ReLU-L2 (5;15) (19;23) (107;3) (108;7)
+SiLU (10;3) (12;8) (17;0) (17;0)
+SiLU-L2 (171;118) (136;89) (17;2) (17;2)
+Softplus (43;14) (65;15) (18;0) (36;14)
+Tanh (37;39) (69;28) (249;126) (272;528)
+Table 3: Mean and standard deviation of the losses of different model setups for the hyperbolic tangent with 77 training points.
+Activation Training MAE $\times 10^{-6}$ Diffusion MSE $\times 10^{-2}$
+$n=7$ $n=8$ $n=7$ $n=8$
+ELU (27;12) (22;32) (41;8) (54;50)
+Leaky ReLU $(\alpha=0.2)$ (7;7) (10;6) (37;38) (54;39)
+LELU $(\beta=0.3)$ (17;14) (57;54) (14;4) (28;10)
+LELU $(\beta=0.4)$ (40;15) (60;36) (8;4) (9;2)
+LELU $(\beta=0.6)$ (35;24) (64;33) (14;6) (16;5)
+ReLU (6;5) (6;4) (37;63) (53;115)
+ReLU-L2 (8;5) (2;2) (72;4) (74;4)
+SiLU (9;5) (10;5) (5;4) (7;3)
+SiLU-L2 (5;106) (170;142) (9;1) (8;0)
+Softplus (21;11) (51;28) (58;35) (71;35)
+Tanh (13;8) (10;22) (96;42) (86;39)
+Tables 2 and 3 present the training and diffusion [^2] In this case for which the function values approach zero, the denominators in Eqs.
+3 and 4 are not implemented.
+losses for $14$ and $7$ training points, respectively.
+As expected, the models tend to perform better with the larger number of training points, however the errors are not necessarily much smaller.
+The models on the lower end of MAEs are generally tending to memorise the training points, as seen in the function plots in Fig.
+5, so the performance of the models is fundamentally gauged by the diffusion metric.
+SiLU and Softplus present adequate diffusion metrics for this hyperbolic tangent function, closely followed by LELU.
+Softplus however loses performance with the deeper model and with the lower number of points.
+All other ac.fs.
+perform rather poorly, and the application of regularisation (strength 1e41e-4) does not improve or even worsens the performance.
+For this particular problem, LELU can perform closely to SiLU at the expense of higher training MAE.
+It is expected that the optimal parameter $\beta$ of LELU is problem and model dependent, which can be observed in the presented results, and that is an intended property.
 (a) 7 training points.
 (b) 14 training points.
-Figure 5: Trained hyperbolic tangent function plots for the different activation functions.
+Figure 5: Trained hyperbolic tangent function plots for the different activation functions for the He-Normal initialisation (no seed).
 4.1.2 Exponential Function
 A 1d exponential equation given by $f(x)=0.1^{x},~x\in[-1,1]$ is used to generate a set of $12$ equally spaced, noiseless training data points.
-The model is trained for 20000 epochs with batch size 3.
-The learning rate is scheduled in the range $[1e-3,2e-7]$ with factor 0.5, patience 500, and cooldown 100.
-Table 3: Losses of the trained models for the exponential function data.
-Activation Training MAE ×$\times 10^{-6}$ Diffusion MSE ×$\times 10^{-3}$
-$n=8$ $n=7$ $n=8$ $n=7$
-Leaky ReLU $(\alpha=0.2)$ 4 31 12 16
-LELU $(\beta=0.3)$ 50 19 3 2
-LELU $(\beta=0.4)$ 24 33 2 2
-LELU $(\beta=0.6)$ 34 62 1 1
-Softplus 120 133 5 2
-SiLU 7 25 13 14
-tanh 21 14 28 37
-ELU 18 12 18 20
-ReLU 4 5 25 19
-Table 3 presents the training losses and diffusion metrics of the trained model.
-As in the previous case, the lower the MAE, the higher is the tendency of the model to memorise the training data, with a resulting increase in the diffusion metric.
-For this case, both SiLU and Softplus show a poorer performance and trade the training and diffusion (overfitting) errors.
-The LELUs are able to adequately meet both low MAE and diffusion errors.
-The trained model predictions are plotted in Fig.
-6.
-As in the previous case, different levels of overfitting are observed with the classical ac.f.’s in this setup, whereas generally robust results are obtained with the LELU.
+The models are trained with $(n\times m)=(7\times 240)$ and $(8\times 240)$, for 15000 epochs with batch size 3.
+The learning rate is scheduled in the range [1e3,2e7][1e-3,2e-7] with factor 0.5, patience 500, and cooldown 100.
+Table 4: Mean and standard deviation of the losses of different model setups for the exponential function with 1212 training points.
+Activation Training MAE $\times 10^{-6}$ Diffusion MSE $\times 10^{-3}$
+$n=7$ $n=8$ $n=7$ $n=8$
+ELU (21;62) (20;15) (12;9) (19;16)
+Leaky ReLU $(\alpha=0.2)$ (10;9) (11;14) (14;4) (15;3)
+LELU $(\beta=0.3)$ (20;6) (36;25) (2;1) (3;2)
+LELU $(\beta=0.4)$ (44;57) (40;99) (2;1) (2;1)
+LELU $(\beta=0.6)$ (38;227) (52;184) (2;0) (1;0)
+ReLU (8;6) (7;8) (21;4) (18;4)
+ReLU-L2 (6;1) (7;2) (18;8) (24;7)
+SiLU (17;95) (12;75) (15;8) (13;7)
+SiLU-L2 (155;683) (104;567) (8;46) (7;5)
+Softplus (115;111) (97;166) (2;1) (4;0)
+Tanh (11;5) (157;165) (33;9) (24;23)
+Table 4 presents the training and diffusion losses of the trained models.
+As in the previous case, the lower the MAE, the higher is the tendency of the model to memorise the training data with a resulting increase in the diffusion metric (see Fig.
+6).
+Different levels of overfitting are observed with the classical ac.fs.
+(including regularisation) in this setup, whereas generally robust results are obtained with LELU.
+Contrary to the previous observations with the hyperbolic tangent results, both SiLU and Softplus show a poorer performance in terms of training MAE or diffusion MSE.
+The LELUs are able to adequately meet both low MAE and diffusion errors, with the tendency for slightly better overall performance with lower $\beta$ due to lower MAE.
 (a) Depth 7.
@@
 The previous problem is reduced to a set of three training points, $x_{i}=(-1,0,1)$, which a single neuron shall learn.
-This exercise can be performed by hand and two of the three points can be exactly matched.
+This simple exercise can be performed by hand so two of the three points can be exactly matched.
 The obtained neuron parameters are used to infer a function as shown in Fig.
@@
 From the perspective of learning nonlinear regression tasks, such weak signal is less then ideal.
-The gradient and Hessian of the loss can be computed by hand for this simple problem.
-This operation shows that the weak ac.f.’s produce a vanishing loss gradient and higher conditioned Hessians, which is indicative of the tendency to poor learning and overfitting.
+The gradient and Hessian of the loss can be computed by hand for this simple problem (motivated by [28]).
+This operation shows that the weak ac.fs.
+produce a vanishing loss gradient and higher conditioned Hessians, which is indicative of the tendency to poor learning and overfitting.
 The LELU on the other hand produces a clear gradient and a lower conditioned Hessian.
 With this simple exercise, the author intends to anecdotally show that the higher “flexibility” of the ac.f.
-is incurring in a weak signal for subsequent neurons and is a potential cause of instability showing as overfitting.
+is incurring in a weak signal for subsequent neurons and is a potential cause of instability showing itself as overfitting.
 4.1.4 Other Findings with the Unidimensional problems
@@
 Figure 8: Trained function plots for the different activation functions on a positive range of $x$.
-It is interesting to note that some ac.f.’s struggle around $x=0$ in both case studies, even when this point is not a training data.
+It is interesting to note that some ac.fs.
+struggle around $x=0$ in both case studies, even when this point is not a training data.
 A thorough explanation is challenging for a large model [24], but the author believes this can be partially understood by a similar analysis of the gradient as in the previous subsection.
@@
 5a, which corroborates the borderline stability of the model.
-Another observation is that discontinuous ac.f.’s generate discontinuous trained models in both exercises.
-In these and some other setups not shown in here, the discontinuous ac.f.’s rather often default to piecewise linear connection between the training points.
-Even when this was not the case, the trained model is still discontinuous, as seen in Figs.
-5, 6 and 8.
-The adopted canonical problems indicate that classical ac.f.’s lack robustness with the problem setup for nonlinear regressions tasks when trained by large models.
+Another observation is that C-1 discontinuous ac.fs.
+generate “polyline” trained models.
+In these and some other setups not shown for conciseness, (leaky) ReLUs rather often default to piecewise linear connection between the training points or at least connect them with several segments of polyline.
+Such patterns are shown in the results in Figs.
+5, 6 and 8, as well for the motor model to be discussed in what follows.
+Interestingly, in none of the cited ReLU-related literature in the introduction, the herein reported “polyline” pattern is commented.
+The adopted canonical problems indicate that classical ac.fs.
+lack robustness with the problem setup for nonlinear regressions tasks when trained by large models.
 The proposition of the objectives in Sect.
 2 for the LELU construction aims at mitigating the reported issues.
-4.1.5 Other Approaches to Decrease Overfitting
-By assuming that the author has been successful in the beginning of this section to justify the most obvious criticism against these exercises, i.e., the adopted model sizes, other strategies also merit further discussions.
-Known strategies to deal with overfitting, namely, regularisation and smaller learning rates for longer epochs were also evaluated.
-The results are not shown for the sake of conciseness since the outcomes are as expected, i.e., a decrease in the overfitting is achieved however at the expense of considerably higher training MAEs and/or computational resources.
-Both L1 and L2 regularisations, with constant strength $41e-$ 4, are able to reduce or remove (L1 regularisation) overfitting, however at the expense of $10x$ or larger training MAEs.
-In terms of learning rates, an example with the ELU required reduction of half of the initial learning rate, and $10x$ longer training epochs.
-The LELU is able to achieve adequate results without those penalties.
-It can robustly learn the nonlinear regression task without memorising the training data.
+LELU is able to achieve adequate results without heavily incurring in such penalties for the presented canonical problems.
+It can robustly learn the nonlinear regression task without memorising the training data, at the expense of slightly higher training MAEs, resembling an inherent but consistent auto-regularisation property of the ac.f.
 This behaviour is further confirmed on the more complex model discussed in what follows.
@@
 A nonlinear feature can also be observed approaching the maximum limits of the $param1$ range.
-The dataset is normalised so as to keep unitary mesh spacing in all dimensions to avoid cell skewness (which also favours the computation of the diffusion differentials), and kept in the positive range to support the more sensitive ac.f.’s.
+The dataset is normalised so as to keep unitary mesh spacing in all dimensions to avoid cell skewness (which also favours the computation of the diffusion differentials), and kept in the positive range to support the more sensitive ac.fs.
 For the current study the trained data must exactly reproduce the noise-free training information [1].
@@
 The model is trained for 9000 epochs with batch size 35.
-The learning rate is scheduled in the range $[1e-3,1e-5]$ ($45e-$ 4 for the models with width $360$ and the $8\times 240$ case) with factor 0.5, patience 250, and cooldown 50.
+The learning rate is scheduled in the range [1e3,1e5][1e-3,1e-5] (5e45e-4 for the models with width $360$ and the $8\times 240$ case) with factor 0.5, patience 250, and cooldown 50.
+ELU, SiLU, and LELU ($\beta=0.4$) are run with 4 seeds of He-Normal and 1 seed of Glorot-Uniform weight initialisers for the $7\times 120$ and $8\times 120$ cases, and the standard deviations are presented as error bars.
+Where applicable, experiments with L2 regularisation are also reported.
 (a) MAE.
 (b) Diffusion SE.
-Figure 10: Trained model losses for the electric motor.
-Figure 10 shows the MAE and diffusion metrics for the different model sizes (indexed by $m^{n}/1e12$ in the plots).
-The tendency to overfit shows as a lack of “convergence” of the MAE with the increase in the model size.
-For this complex problem the mean errors do not show large differences between the “good” and “bad,” so visual inspection is also necessary.
-Nonetheless, it can be seen that both metrics are lower with the LELU and are also more robust to model size.
-SiLU and Softplus do not present any favourable behaviour as partially seen with the canonical problems.
-The conclusions above are corroborated by the slice plots in Figs.
+Figure 10: Electric motor model losses for different ac.fs. and model sizes indexed by $m^{n}/1e12$.
+Figure 10 shows the MAE and diffusion metrics for the different model sizes and ac.fs.
+The tendency to lose generalisation shows as a lack of “convergence” of the losses with the increase in model size.
+For this complex problem the mean errors do not show large differences between the “good” and “bad,” so visual inspection guided by the results of the diffusion loss is necessary.
+In overall, similar behaviours as with the canonical problems are also observed in this case.
+In special, the most promising classical ac.fs.
+SiLU and Softplus also present poor performance, consistent with the slice plots in Figs.
 11 and 12.
+LELU successfully achieves lower losses amongst the tested ac.fs.
+and is also more robust to model size.
+As expected, the losses with different values of $\beta$ can be sensitive to model parameters.
+As an example, $\beta=0.2$ presents better results for smaller models and loses performance with larger ones, whereas $\beta=0.6$ presents the opposite trend.
+LELU with $\beta=0.4$ presents almost iso-performance with model parameters, and the sensitivity (error bars) to initialisation is also the smallest.
+In all cases, the results are more promising than any other ac.f.
 (a) LELU $(\beta=0.2)$
@@
 Figure 11 shows slices along $param2$, which is the fast-changing variable.
-The tanh and ReLU results are not shown due to the poor performance.
-The region $param2=[0,2]$, where the highest slopes are seen, presents a tough task for the ac.f.’s.
+The tanh and ReLU results are not shown due to poor performance.
+The region $param2=[0,2]$, where the highest slopes are seen, presents a tough task for the ac.fs.
 Considerable “localised” overfitting is observed for the ELU, SiLU and Softplus.
-Interestingly, the less flexible leaky ac.f.’s show lower overfitting content and also less dependency on the model size.
+Interestingly, the less flexible leaky ac.fs.
+show lower overfitting content and also less dependency on the model size.
 This nice behaviour prompted the current effort into assembling this property into a new ac.f.
@@
 with different model sizes and the capacity to code small values.
-The less flexible leaky ac.f.’s clearly show tighter and more representative capacity.
+The less flexible leaky ac.fs.
+clearly show tighter and more representative capacity.
 SiLU and Softplus are also able to demonstrate a tighter spread, albeit with an offset error to the training data, whereas ELU is generally poorly behaved.
@@
 1, this ac.f.
-slowly changes slopes and is therefore smoother than other ac.f.’s, thereby enabling a stronger a priori fitness for generalisable nonlinear regressions.
+slowly changes slopes and is therefore smoother than other ac.fs., thereby enabling a stronger a priori fitness for generalisable nonlinear regressions.
 (a) $6\times 120$
@@
 Figure 13: Model size effect on the electric motor predictions for different activation functions.
-For this motor model the LELU with $\beta=0.4$ provides the most agreeable results.
 Upon careful visual inspection of each NN result (see Fig.
@@
 Figure 14: Regularisation effect on the electric motor predictions for the $8\times 120$ model.
-It is worth revisiting the discussion in Sect.
-4.1.5.
-Except from reducing the model size, which is in fact hopefully clear now the need for such larger models, the other techniques to control overfitting also apply to the motor model.
-As shown in Fig.
-14, both L1 and L2 regularisations (strength $41e-$ 4) reduce the overfitting behaviour of the SiLU, at the expense of representation of training points.
-The MAE increase by more than a factor of 6 in these cases and the performance does not improve over the LELU ($\beta=0.4$).
+Figure 14 shows slices with L2 (strength 1e41e-4) regularisation to reduce the overfitting tendency of ReLU and SiLU.
+As corroborated by the results in Fig.
+10, the diffusion loss indicates a reduction in overfitting at the expense of an increase in the training MAE by more than a factor of 6.
+The performance does not improve over the LELU ($\beta=0.4$) and ReLU still keeps its “kinky” piecewise-linear predictions.
 Another strategy is to implement a change of variables, for instance a power correction $\tilde{y}=y^{0.1}$, to reduce the large differences in the max and min values.
@@
 and is a trainable parameter.
-Available nonlinear ac.f.’s have generally high flexibility scores which seem to be associated with proneness to overfitting.
+Available nonlinear ac.fs.
+have generally high flexibility scores which seem to be associated with proneness to overfitting.
 A null flexibility score can only represent linear data, on the other hand.
 Smoothness is sought to avoid coding discontinuities at the predictions of the trained model.
-This is shown to be a result of the ReLU and Leaky ReLU ac.f.’s.
+This is shown to be a result of the ReLU and Leaky ReLU ac.fs.
 The Leaky ReLU however has a trainable flexibility and this is shown to support the claim of beneficial and robust behaviour of this property for nonlinear regression tasks.
@@
 Its properties implicitly produce a behaviour similar to regularisation without strong penalisation of the training errors.
-The LELU does not show evidence of trading off training error with the overfitting/testing metric, at least not as strongly as other ac.f.’s.
-It also shows less sensitivity with model size and training points for nonlinear regression tasks.
+The LELU does not show evidence of trading off training error with the overfitting/testing metric, at least not as strongly as other ac.fs.
+It also shows less sensitivity with model size and training sample size for nonlinear regression tasks.
+The parameter $\beta$ is proposed so it can be dedicatedly trained for better generalisation for different problems.
 Another proposition in the current work is a metric to identify overfitting.
@@
 The author expects the proposed methods to be generalisable to other machine learning or dataset strategies.
-The observed properties of the proposed LELU should reduce the workload to select a good model setup and is as a such a highly desirable behaviour.
+The observed properties of the proposed LELU should reduce the workload to select a good model setup and is as such a highly desirable behaviour.
 The Laplace operator with its measure of the diffusion of a system provides a powerful tool for detection (and prevention) of overfitting in data regression.